### v1

# Vectorized Computation of Euler Characteristic Functions and Transforms

Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are neither optimized for speed nor do they scale to higher-dimensional settings. In this work, we present a vectorized framework for computing such topological transforms using tensor operations, which is highly optimized for GPU architectures and works in full generality across geometric simplicial complexes (or cubical complexes) of arbitrary dimension. Experimentally, the framework demonstrates significant speedups (up to $180\times$) over existing methods when computing the WECT and ECF across a variety of image datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT. [^1] The pyECT package is available at https://github.com/compTAG/pyECT
1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant: the Euler characteristic [euler1758elementa]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [edel2002topological] (resulting from applying homology to filtrations) and Euler characteristic functions [Serra1982] (resulting from applying the Euler characteristic to filtrations).
Topological transforms [Turner2014PHT] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a height filtration of $K$ is constructed and then a topological invariant is applied to each height filtration. Applying the Euler characteristic (EC) to each height filtration results in the Euler characteristic transform (ECT) [Turner2014PHT, munch2023invitation], the focus of this paper. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [curry2022many, Ghrist2018]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
A useful variant of the ECT is the weighted Euler characteristic transform (WECT) [Jiang2020TheWE]. The WECT extends the ECT to the setting of weighted geometric complexes in $\mathbb{R}^{n}$. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [Jiang2020TheWE]. Moreover, the WECT has been employed in a broad range of applications. In the field of oncology, the WECT has been used to predict survival rates of patients with glioblastoma multiforme brain tumors [crawford2020predicting]. In plant biology, the WECT has been used to classify seed phenotypes [amezquita2022measuring]. The WECT has also proven particularly adept at image classification tasks [cisewski2023weighted, Jiang2020TheWE].
Numerous software packages already exist for computing the WECT, including Demeter [amezquita2022measuring], SINATRA [crawford2020predicting], EuCalc [lebovici2024efficientcomputationtopologicalintegral], and FastTopology [laky2024]. The packages Demeter and SINATRA processes arbitrary geometric simplicial and cubical complexes in Euclidean spaces, but are not optimized for speed. The packages FastTopology and EuCalc are optimized for speed, but only apply to axis-aligned cubical complexes (i.e., grids) in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. Two other packages, [roell2023differentiable] and [saxena2025scalable], provide vectorized implementations optimal for Graphics Processing Unit (GPU) hardware. However, these packages compute an approximation of the ECT known as the differentiable Euler characteristic transform (DECT). Furthermore, [saxena2025scalable] is an optimization for the specific case of unweighted cubical complexes in $\mathbb{R}^{2}$ or $\mathbb{R}^{3}$. As outlined in [roell2023differentiable], the DECT provides several advantages for computation in neural network applications. However, our method is focused on exact computation of the WECT in full generality.
Contributions
In this work, we provide a fully vectorized framework for computing WECFs and WECTs. Our implementation applies in full generality to arbitrary geometric simplicial and cubical complexes in Euclidean spaces, and supports automatic gradient computations. The vectorized operations are highly optimized for fast computation on modern GPU hardware. Moreover, our implementation significantly outperforms existing software packages, including the optimized packages EuCalc and FastTopology.
2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to simplicial complexes in this paper; however, our results and code work for cubical complexes with no modifications necessary.
2.1 Lower-Star Filtrations
While the primary focus of this paper is on the WECT, our algorithm applies more generally to any finite set of lower-star filtrations. Here, we provide the necessary background on lower-star filtrations.
Definition 2.1 (Lower-Star Filtrations).
Let $K$ be a simplicial complex with vertex set $K_{0}$ and let $\operatorname{sub}(K)$ denote the set of subcomplexes of $K$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f:K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function [^2] In this paper, we use a lower case letter to denote a filter function and the same upper case letter to denote the filtration it induces.
$F:\mathbb{R}\to\operatorname{sub}(K),$
defined as
$F(t):=\{\sigma\in K\mid\max_{v\in\sigma}f(v)\leq t\}$ (1)
where the max is taken over all vertices $v$ of the simplex $\sigma$.
The main examples of lower-star filtrations used in this paper are height filtrations. To define height filtrations, we first define geometric simplicial complexes. We adopt the definition from [edelsbrunner2010computational, Chap. 3, pp. 52], and refer the reader to this source for further details.
Definition 2.2 (Geometric Simplicial Complex).
A geometric simplicial complex is a finite collection of geometric simplices $K$ in $\mathbb{R}^{n}$ that satisfies the following two conditions:
  • If $\sigma\in K$, then all faces of $\sigma$ are in $K$.
  • If $\sigma,\tau\in K$, then $\sigma\cap\tau$ is either empty, or a face of both $\sigma$ and $\tau$.
Note that any (abstract) simplicial complex has a (non-unique) geometric realization in Euclidean space, as demonstrated, for example, in [edelsbrunner2010computational, Chap. 3, pp. 53]. Given an abstract simplicial complex $K$ with underlying space $|K|$, a function $\ell:|K|\to\mathbb{R}^{n}$ that provides a geometric realization of $K$ in $\mathbb{R}^{n}$ is called an embedding of $K$ into $\mathbb{R}^{n}$.
When working with geometric realizations of abstract simplicial complexes, we often abuse notation by avoiding explicit reference to the embedding function $\ell$. Consequently, when we refer to a vertex $v\in K_{0}$, we often mean its image $\ell(v)\in\mathbb{R}^{n}$.
Definition 2.3 (Height Functions and Filtrations).
Let $K$ be a simplicial complex embedded in $\mathbb{R}^{n}$ via the map $\ell:|K|\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{\mathbb{S}}^{n-1}$} be a direction vector. The height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}:K_{0}\to\mathbb{R}$} defined as
$h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$
The height filtration of $K$ in direction $\mathbf{s}$, which we denote by $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
2.2 The Weighted Euler Characteristic Transform
Here, we introduce the main focus of this paper: the WECT.
Definition 2.4 (Weighted Simplicial Complexes).
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega:K\to\mathbb{R}$ is a function (referred to as the weight function).
Weighted simplicial complexes are often used to model various types of data. For example, triangular meshes are widely used to model three-dimensional surfaces, with weight functions capturing scalar fields on the surface. Another example comes from image data, where grayscale images are converted into weighted simplicial complexes via the Freudenthal triangulation [cisewski2023weighted, Jiang2020TheWE] with weights derived from the pixel intensities.
Originally introduced by Euler in 1758 for convex polyhedra [euler1758elementa], the Euler characteristic (EC) is a topological invariant that readily extends to simplicial complexes. We use an extension of the EC to weighted simplicial complexes: the weighted Euler characteristic.
Definition 2.5 (Weighted Euler Characteristic).
The weighted Euler characteristic (WEC) of a weighted simplicial complex $(K,\omega)$ is
$\chi(K,\omega):=\sum_{\sigma\in K}(-1)^{\dim{\sigma}}\omega(\sigma).$
Remark 2.1.
The Euler characteristic is a special case of the WEC. For a simplicial complex $K$, its Euler characteristic is the WEC of $(K,\mathbf{1})$, where $\mathbf{1}:K\to\mathbb{R}$ is the constant function that assigns value one to each simplex.
While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex $K$. To incorporate more information, we apply the WEC to a lower-star filtration of $K$, resulting in a weighted Euler characteristic function (WECF).
Definition 2.6 (Weighted Euler Characteristic Function).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $f\colon K_{0}\to\mathbb{R}$ be a vertex filter on $K$. The weighted Euler characteristic function  [^3] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $f$ is the function
$\operatorname{WECF}_{f}:\mathbb{R}\to\mathbb{R}$
defined as
$\operatorname{WECF}_{f}(t):=\chi\big(F(t),\omega|_{F(t)}\big)$ (2)
where $F$ is the lower-star filtration of $K$ induced by $f$.
Definition 2.7 (Directional WECF).
Let $\mathbb{K}=(K,\omega)$ be a weighted geometric simplicial complex in $\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The directional WECF of $\mathbb{K}$ in direction $\mathbf{s}$ is the WECF of $\mathbb{K}$ induced by the height function $h_{\mathbf{s}}$.
Remark 2.2.
In image analysis tasks (and their three-dimensional analogs), there are two common choices of vertex filters considered. The first uses the pixel intensities (i.e., grayscale values) as the vertex filter. The implementations in [laky2024, wang2022gpu] compute the result of applying the (unweighted) Euler characteristic to the corresponding lower-star filtration. We refer to this simply as the ECF of an image. The second uses the pixel intensities as weights, and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [amezquita2022measuring, tang2022topological, lebovici2024efficientcomputationtopologicalintegral].
Example 2.1.
An illustration of the directional ECF and directional WECF is presented in Figure 1. The full simplicial complex, including weighted vertices, is displayed in Figure 1(c). The filtration is along direction $(1,0)$, and the corresponding ECF and WECF are included in Figure 1(d) and Figure 1(e), respectively. The ECF and WECF are the same until time 1 when the weights of the two new vertices have weights of 0.5.
(a) Time=0
(b) Time=1
(c) Time=2
(d) ECF in direction (1,0)
(e) WECF in direction (1,0)
Figure 1: Filtration of a simplicial complex (a-c) along direction (1,0), where the weights are specified in the green zero-simplices. The weight assigned to a simplex $\sigma$ is the maximum weight of its zero-simplex faces. The corresponding ECF (d) and WECF (e) are displayed in the direction (1,0); see Equation (2). At times 1 and 2, the ECF and WECF differ due to the WECF’s use of the weighted simplices.
Typically, there is no canonical direction vector to choose for directional WECFs. For this reason, it is common to choose all direction vectors. This results in the weighted Euler characteristic transform (WECT).
Definition 2.8 (Weighted Euler Characteristic Transform).
Let $\mathbb{K}=(K,\omega)$ be a weighted geometric simplicial complex in $\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function
$\operatorname{WECT}_{\mathbb{K}}:\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$
defined as
$\operatorname{WECT}_{\mathbb{K}}(\mathbf{s},t):=\operatorname{WECF}_{h_{\mathbf{s}}}(t)=\chi\big(H_{\mathbf{s}}(t),\omega\big).$ (3)
2.3 Existing Frameworks
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings. Definition 2.8 suggests a naïve algorithm for computing a discretization of the WECT:
  1. 1.
    Choose a finite set of directions $S=\{s_{1},\ldots,s_{m}\}\subsetneq\mathbb{S}^{n-1}$ and a finite set of height thresholds $T=\{t_{1},\ldots t_{\ell}\}\subsetneq\mathbb{R}$.
  2. 2.
    Construct the simplicial complexes $\{\sigma\mid 0pt{\sigma}{s}\leq t\}$ for each $s\in S$ and $t\in T$.
  3. 3.
    Compute the WECT of each complex.
Suppose that $n=|K|$ is the number of simplices in $K$. This naïve implementation runs in $\Theta(\ell mn)$ time, This can be improved to $\Theta(m(\ell+n\log n))$ time using a sorting-based approach. There are several publicly available implementations to compute the WECT and its variants, which are described next.
Demeter
Demeter [amezquita2022measuring] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT.
SINATRA
The SINATRA package [tang2022topological] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary simplicial complexes. Note that it does not allow for weighted complexes and is not optimized for speed.
EuCalc
EuCalc [lebovici2024efficientcomputationtopologicalintegral] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes.
FastTopology
The FastTopology framework [laky2024] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF.
GPU ECC
GPU ECC [wang2022gpu] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems.
In this paper, we present the pyECT framework for computing the WECT. In general, the pyECT framework has several advantages over these existing implementations. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across axis-aligned cubical complexes, arbitrary cubical complexes, and geometric simplicial complexes; see Table 1. Second, pyECT provides frameworks for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the vectorized operations used in pyECT are highly optimized for modern GPU hardware, which, in most cases, results in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the simplicity of the vectorized approach makes it very advantageous for software implementation, especially in the context of well-maintained, abstracted tensor computation libraries like PyTorch [pytorch1, ansel2024pytorch]. For example, the core WECT algorithm in the pyECT implementation is only about $25$ lines of code.
One minor disadvantage of pyECT is that it does not compute exact Euler characteristic values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT because directional WECFs do not have consistent heights at which they change values across varying directions. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [Jiang2020TheWE, Turner2014PHT, munch2023invitation, amezquita2022measuring]). While it is possible to compute a vectorization from exact values using EuCalc, it requires an additional computation step that is linear in the size of the discretization.
Table 1: Comparison of acceptable input types of ECT packages.
Package
Simplicial
Complexes
Cubical
Complexes
Axis-Aligned
Cubical Complexes
pyECT yes yes yes
EuCalc no no yes
FastTopology no no yes
Demeter no no yes
SINATRA yes yes yes
2.4 Tensor Operations
Our algorithms almost exclusively use vectorized operations on tensors. These have two main advantages over non-vectorized approaches. The first advantage is that the vectorized operations we use have highly optimized implementations in PyTorch [pytorch1, ansel2024pytorch] that run in parallel on both CPUs and GPUs. The second advantage is that they support automatic gradient computations, facilitating the use of our implementation in neural networks. Here, we review tensors and the vectorized operations used in our algorithm.
For any positive integer $m$, let $m]]:=\{0,1,\ldots,m-1\}$. For $\{m_{i}\}_{i=0}^{d-1}\subset\mathbb{Z}_{+}$, a $d$-dimensional tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ is a function
$T\colon[[m_{0}]]\times[[m_{1}]]\times\cdots\times[[m_{d-1}]]\to\mathbb{R}.$
We exclusively use one-dimensional tensors (i.e., vectors), two-dimensional tensors (i.e., matrices), and three-dimensional tensors in our algorithm. A specific dimension of a tensor is called an axis.
Next, we introduce the tensor operations utilized in our algorithm. For examples of the following tensor operations, we refer the reader to Appendix A.
  1. 1.
    Zeros: Given a shape $L$, $\texttt{zeros}(L)$ is the tensor with shape $L$, where every value is 0.
  2. 2.
    Element-wise operations: Any function $f:\mathbb{R}\to\mathbb{R}$ can be applied to tensors element-wise by post-composition. That is, if $T$ is a tensor, then $f(T)$ is the tensor with the same shape as $T$ and values
    $f(T)[i_{0},i_{1},\ldots,i_{d-1}]=f\big(T[i_{0},i_{1},\ldots,i_{d-1}]\big).$
  3. 3.
    Matrix multiplication: If $T$ and $S$ are tensors with shape $(m_{0},m_{1})$ and $(m_{1},m_{2})$ respectively, we use the notation TST*S to denote their matrix multiplication.
  4. 4.
    Cumulative sum: Let $T$ be a tensor of shape $(m_{0},m_{1})$. Then, $\texttt{CumulativeSum}(T)$ is the tensor with shape $(m_{0},m_{1})$ and values
    $\texttt{CumulativeSum}(T)[i,j]=\sum_{\begin{subarray}{c}k\in[[m_{1}]]\\ k\leq j\end{subarray}}T[i,k].$
    In this work, we only use the cumulative sum on two-dimensional tensors over the second axis, $m_{1}]]$, as defined here. In more general settings, the cumulative sum can be defined over any axis of a tensor.
  5. 5.
    Reduce maximum: Let $T$ be a tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ and let $p\in[[d]]$. Then, $\texttt{ReduceMax}(T,p)$ is the tensor with shape
    $(m_{0},\ldots,m_{p-1},m_{p+1},\ldots,m_{d-1})$
    and values
    $\texttt{ReduceMax}(T,p)[i_{0},\ldots,i_{p-1},i_{p+1},\ldots i_{d-1}]=\max_{j\in[[m_{p}]]}T[i_{0},\ldots,i_{p-1},j,i_{p+1},\ldots i_{d-1}].$
  6. 6.
    Advanced indexing: Let $T$ be a tensor with shape $(m_{2},m_{3})$, and let $I$ be a tensor with shape $(m_{0},m_{1})$ and values in $m_{2}]]$. The tensor $T[I]$ has shape $(m_{0},m_{1},m_{3})$ and values
    $T[I][i,j,k]=T\big[I[i,j],k\big].$
  7. 7.
    Scatter add: Let $S$ be a tensor with shape $(m_{0},m_{1})$, let $I$ be a tensor with shape $(m_{0},m_{2})$ and values in $m_{1}]]$, and let $T$ be a tensor with shape $(m_{2})$. The function $\texttt{ScatterAdd}(I,T)$ updates $S$ in-place by adding to it, the difference tensor $D$ with shape $(m_{0},m_{1})$ and values
    $D[i,j]=\sum_{\begin{subarray}{c}k\in[[m_{2}]]\\ I[i,k]=j\end{subarray}}T[k].$ (4)
    That is, for a tensor $S$, we have
    $\texttt{ScatterAdd}(I,T)(S)=S+D,$
    where ++ is taken element-wise and $S$ is updated in place. In practice, the tensor $D$ is not explicitly built, but is helpful for analysis.
3 Computation of WECFs
In this section, we present a new algorithm for computing the WECFs of a weighted simplicial complex with respect to a finite set of vertex filters. Computation of the WECT of a weighted geometric simplicial complex is the motivation for Algorithm 1, and corresponds to the case where the vertex filters are height functions. See Example 3.1 for specifics on how Algorithm 1 may be used to compute the WECT.
3.1 Algorithm
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m}\}$ for the entirety of this section. For any $i\geq 0$, we denote by $K_{i}$, the set of $i$-simplices in $K$ and we let $k_{i}=|K_{i}|$. We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}}^{i}\}$ for each $i\geq 1$. To simplify notation, we extend the vertex filters $f_{p}:K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}:K\to\mathbb{R}$ as
$f_{p}^{\max}(\sigma)=\max_{v\in\sigma}f_{p}(v).$
Our algorithm uses the following inputs
  • Complex: A list of tuples:
    • $\texttt{Complex}[0]=(\texttt{Filters},\texttt{VertexWeights})$
        Filters is a tensor with shape $(k_{0},m)$ and values
        $\texttt{Filters}[a,p]=f_{p}(v_{a}).$
        VertexWeights is a tensor with shape $(k_{0})$ and values
        $\texttt{VertexWeights}[a]=\omega(v_{a}).$
    • for $i\geq 1$: $\texttt{Complex}[i]=(i\texttt{-SimplexVertices},i\texttt{-SimplexWeights})$
        $i\texttt{-SimplexVertices}$ is a tensor with shape $(k_{i},i+1)$ and values in $k_{0}]]$. The rows of this tensor correspond to the vertex sets of each $i$-simplex.
        $i\texttt{-SimplexWeights}$ is a tensor with shape $(k_{i})$ and values
        $i\texttt{-SimplexWeights}[b]=\omega(\sigma_{b}^{i}).$
  • NumVals: The number of values to sample the filter functions over.
Define MaxVal as
$\texttt{MaxVal}=\max_{\begin{subarray}{c}p\in[[m]]\\ v\in K_{0}\end{subarray}}|f_{p}(v)|.$
We discretize $\texttt{MaxVal},\texttt{MaxVal}]$ by evenly sampling NumVals points. That is, we define the map
$\beta\colon[[\texttt{NumVals}]]\to[-\texttt{MaxVal},\texttt{MaxVal}]$
by
$\beta(q)=q\frac{2*\texttt{MaxVal}}{\texttt{NumVals}-1}-\texttt{MaxVal}.$
The function $\beta$ has a left inverse
$\alpha\colon[-\texttt{MaxVal},\texttt{MaxVal}]\to[[\texttt{NumVals}]]$
defined as
$\alpha(t)=\bigg\lceil\frac{(\texttt{NumVals}-1)(\texttt{MaxVal}+t)}{2*\texttt{MaxVal}}\bigg\rceil.$ (5)
The maps $\alpha$ and $\beta$ are order-preserving and satisfy
$\alpha(t)\leq q\iff t\leq\beta(q)$ (6)
for all $t\in[-\texttt{MaxVal},\texttt{MaxVal}]$ and $q\in[[\texttt{NumVals}]]$. In other words, $\alpha$ and $\beta$ form a Galois connection.
[…]
0:Complex, NumVals
0:WECFs: tensor of shape $(m,\texttt{NumVals})$ and values $\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}(\beta(q))$.
[…]
5:for $i=1$ to $\dim(K)$ do
6:  SimplexVertices, SimplexWeights $\leftarrow\texttt{Complex}[i]$
7:  SimpIndices $\leftarrow\texttt{VIndices}[\texttt{SimplexVertices}]$
8:  $\texttt{MaxSimpIndices}\leftarrow\texttt{ReduceMax}(\texttt{SimpIndices},1)$
9:  $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})(\texttt{diffWECFs})$
10:end for
[…]
12:
13:return WECFs
Algorithm 1 ultimately produces a tensor WECFs with shape $(m,\texttt{NumVals})$ and values
$\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$
The algorithm begins by initializing a zero-tensor diffWECFs with shape $(m,\texttt{NumVals})$. Next, the vertex filter values are mapped to indices in $\texttt{NumVals}]]$ by applying the map $\alpha$ elementwise to the tensor Filters. This gives the tensor VIndices with values
$\texttt{VIndices}[a,p]=\alpha\big(f_{p}(v_{a})\big).$ (7)
The operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is then applied to diffWECFs, resulting in adding to diffWECFs, the tensor $D_{0}$ with shape $(m,\texttt{NumVals})$ and values
$D_{0}[p,q]=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]=q\end{subarray}}\texttt{VertexWeights}[a].$
Next, inside the for-loop, fix an $i\in\{1,2,\ldots\dim(K)\}$. The tensor SimplexVertices has rows indexed by $K_{i}$ with the entries in each row corresponding to the vertex sets of simplices in $K_{i}$. Therefore, the tensor MaxSimpIndices has shape $(k_{i},m)$ and values
$\texttt{MaxSimpIndices}[b,p]$ $=\max_{j\in[[i+1]]}\texttt{VIndices}\big[\texttt{SimplexVertices}[b,j],p\big]$
$=\max_{v\in\sigma_{b}^{i}}\alpha\big(f_{p}(v)\big)$
$=\alpha\big(\max_{v\in\sigma_{b}^{i}}f_{p}(v)\big)\hskip 42.67912pt\text{as }\alpha\text{ is order-preserving}$
$=\alpha\big(f_{p}^{\max}(\sigma_{b}^{i})\big).$ (8)
The operation $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$ updates diffWECFs by adding to it, the tensor $D_{i}$ with shape $(m,\texttt{NumVals})$ and values
$D_{i}[p,q]=\sum_{\begin{subarray}{c}b\in[[k_{i}]]\\ \texttt{MaxSimpIndices}[b,p]=q\end{subarray}}(-1)^{i}\texttt{SimplexWeights}[b].$
After the conclusion of the for loop, we have
$\texttt{diffWECFs}=\sum_{i=0}^{\dim(K)}D_{i}.$
Finally, a cumulative sum is applied to diffWECFs to obtain the tensor WECFs. From the fact that the cumulative sum is a linear operator, we have
$\texttt{WECFs}=\sum_{i=0}^{\dim(K)}\texttt{CumulativeSum}(D_{i}).$
The cumulative sum of $D_{0}$ has values
$\texttt{CumulativeSum}(D_{0})[p,q]$ $=\sum_{q^{\prime}\leq q}\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]=q^{\prime}\end{subarray}}\texttt{VertexWeights}[a]$
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]\leq q\end{subarray}}\texttt{VertexWeights}[a]$
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \alpha(f_{p}(v_{a}))\leq q\end{subarray}}\texttt{VertexWeights}[a]$ by Equation (7)
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ f_{p}(v_{a})\leq\beta(q)\end{subarray}}\texttt{VertexWeights}[a]$ by Equation (6)
$=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v).$ (9)
Likewise, for any $i\in\{1,\ldots,\dim(K)\}$, Equation (3.1) together with the same argument used to show Equation (3.1) gives
$\texttt{CumulativeSum}(D_{i})[p,q]=\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma).$
Therefore, the tensor WECFs has values
$\texttt{WECFs}[p,q]$ $=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v)+\sum_{i=1}^{\dim(K)}\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
$=\sum_{\begin{subarray}{c}\sigma\in K\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
$=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$
Example 3.1.
Computing a discretization of the WECT of a weighted geometric simplicial complex $\mathbb{K}=(K,\omega)$} in $\mathbb{R}^{n}$ is a special case of Algorithm 1 as follows. Consider a tensor $D$, with shape $(d,n)$, of direction vectors over which to sample the WECT. For any $p\in[[d]]$, let $\mathbf{s}_{p}\in\mathbb{R}^{n}$ be the $p$-th row of $D$. Let $V$ be the tensor, with shape $(k_{0},n)$, of the vertex coordinates of $K$. Then, the tensor $V*D^{T}$ has values
$(V*D^{T})[a,p]=h_{\mathbf{s}_{p}}(v_{a}).$
Therefore, using the tensor $V*D^{T}$ as the Filters tensor in Algorithm 1 produces a discretization of the WECT over the finite sample of directions in $D$.
3.2 Complexity
Here, we analyze the asymptotic complexity of Algorithm 1. We make the assumption that the ambient dimension $n$ is fixed. Let $k$ be the total number of simplices of $K$, let $m$ be the number of vertex filters, and let NumVals be the number of values to sample the vertex filters over.
The function ComputeWECFs has inputs Complex and NumVals requiring $\Theta(km)$ memory usage. The algorithm first initializes a zero tensor diffWECFs of shape $(m,\texttt{NumVals})$, requiring $\Theta(m\texttt{NumVals})$ memory. The function $\alpha$ (see Equation (5)) is then applied elementwise to Filters, taking $\Theta(km)$ time.
Next, the operation
$\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$
is applied to diffWECFs. ScatterAdd adds a value from VertexWeights to diffWECFs for each entry of $\texttt{VIndices}^{T}$, resulting in a time complexity of $\Theta(km)$ and no additional memory usage as the operation is performed in place.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
with shape $(k_{i},i+1,m)$ is constructed, taking $\Theta(km)$ time and memory. This tensor is then reduced along the second axis by taking the maximum values, creating the tensor MaxSimpIndices with shape $(k_{i},m)$ in $\Theta(km)$ time and memory. The tensor diffWECFs is then updated by applying the map
$\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$
which takes $\Theta(km)$ time and memory.
After the conclusion of the for-loop, we have a total time and memory complexity of
$\Theta\big(nkm\big)=\Theta(km).$
Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in a final time and memory complexity of $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1.
In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a theoretically optimal final time and memory complexity of $\Theta(k)$ for Algorithm 1.
4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods on a variety of datasets.
4.1 Experimental Design
The following section outlines the experimental design, including the datasets used, the algorithms compared, and the parameters used in the experiments. Experiments were performed on a single node with a single CPU or a single GPU (where applicable). The CPU used was an EPYC Rome, and the GPU used was an NVIDIA A40. Run times were recorded using the timeit Python library with GPU synchronization (to prevent inaccurate timing from asynchronous GPU operations).
Datasets
The methods were evaluated on both image and three-dimensional point cloud datasets. In particular, we use:
  • Fashion-MNIST [xiaoFashionMNIST]. Fashion-MNIST is a dataset consisting of $28\times 28$ grayscale images of clothing items. In our experiments we use the training set, which consists of $60,000$ images.
  • ImageNet [deng2009ImageNet]. The Large Scale Visual Recognition Challenge (ILSVRC) 2012-2017 subset of ImageNet was used. From the $50,000$ validation images, we randomly selected $1,000$ images with a dimension (horizontal or vertical) of between $200$ and $500$ pixels. This creates a dataset of images that generate complexes large enough to demonstrate the computational efficiency of our method, while still being small enough to avoid memory constraints in our computation environment.
Implementations
Section 2.3 outlines a variety of existing methods for computing topological functions and transforms. As discussed in Remark 2.2, methods like EuCalc, Demeter, and SINATRA compute the directional (W)ECF and (W)ECT, while methods like FastTopology and GPU ECC compute the ECF. For computation of the WECT, we compare our method with EuCalc  [^4] The implementation of EuCalc was accessed at https://github.com/HugoPasse/Eucalc. . Note that Demeter and SINATRA are not optimized for speed, and are thus not included in the comparison. Furthermore, Demeter is included in a comparison with EuCalc in [lebovici2024efficientcomputationtopologicalintegral], where the implementation is shown to be significantly slower than EuCalc. In the case of computing the ECF, we compare our method with FastTopology  [^5] The implementation of FastTopology was accessed at https://github.com/zavalab/ML/tree/master/FastTopology. . Note that the implementation of GPU ECC is only available on Windows operating systems, which was not available in our computation environment.
4.2 Results
We now present the experimental results, measured by the speedup factor, of our method over the existing methods. Speedup factor is defined as the ratio of the average runtime of the existing baseline method (either EuCalc or FastTopology) divided by the average runtime of the proposed method (pyECT). The first section explores the computation of the WECT, while the second section explores the computation of the ECF.
4.2.1 WECT Computation
In this section, we present the results of computation of the WECT using our vectorized method. We provide a comparison of the runtime of our method with EuCalc on the Fashion-MNIST and ImageNet datasets. While pyECT significantly outperforms EuCalc on the Fashion-MNIST dataset, the speedup is far more pronounced on the ImageNet dataset. This is because Fashion-MNIST consists of $28\times 28$ pixel images while the images we selected from ImageNet are significantly bigger.
The results for Fashion-MNIST are shown in Figure 2. As the number of directions increases, the speedup of our method over EuCalc increases linearly. We note that increasing the number of heights used in the vectorization of our method (which is not a parameter of EuCalc) does not have a significant effect on the speedup. At $25$ directions, our method is approximately twice as fast as EuCalc.
Figure 2: Speedup of pyECT over EuCalc on WECT computation of Fashion-MNIST.
The results for ImageNet are shown in Figure 3. As with Fashion-MNIST, as the number of directions increases, the speedup of our method over EuCalc increases. However, as the complexes are significantly larger, for $25$ directions the speedup of our method ranges from $25$ times to $170$ times. As EuCalc does not explicitly depend on the number of heights, its runtime remained constant as the number of heights used increases. However, our method’s speedup factor increases as the number of heights increases (despite requiring more vector computation). While this appears counterintuitive, in reality the vectorized code encounters less efficient memory access patterns at lower numbers of heights, which leads to a smaller advantage over EuCalc with the smaller tensor values. For a technical description of memory access patterns specific to the GPU hardware used in these experiments, we refer the reader to [gpu2025].
Figure 3: Speedup of pyECT over EuCalc on WECT computation of ImageNet.
For both of these datasets, pyECT demonstrates a significant speedup over EuCalc. Note, however, that EuCalc computes exact height values for the WECT, whereas our method computes a vectorization of the WECT. Of course, in many applied settings, vectorizations are more desirable than exact heights due to their utility in machine learning and statistical pipelines. Obtaining a vectorization from EuCalc involves an additional step that further slows it down in comparison to pyECT. The above experiments demonstrate that pyECT admits large speedups when compared to EuCalc, even without the additional step of producing a vectorization in EuCalc.
4.2.2 ECF Computation
In this section, we present the results of computation of the ECF using our vectorized method. We compare our method with FastTopology on the Fashion-MNIST and ImageNet datasets. Additionally, a “padded” ImageNet dataset was added, which increased the dimensions of each image to a size of $1000\times 1000$ pixels, with intensities drawn uniformly at random. The padded dataset was added because for smaller datasets (e.g., Fashion-MNIST), the overhead of parallelization can mitigate the computational speedups due to the smaller size of the complexes. For pyECT, the number of heights was fixed to $256$ to match that of FastTopology (and compute across each intensity). The results are presented in Figure 4.
Across all datasets, pyECT demonstrates a significant speedup over FastTopology. For the single core case, due to the small size of the complexes in Fashion-MNIST, the results of the two methods are comparable. However, in the single core case for the larger datasets, pyECT demonstrates a speedup of nearly $20$ times for the unpadded dataset, and nearly $30$ times for the padded dataset. As the number of cores used in FastTopology increased, the advantage of pyECT decreased slightly. However, pyECT still demonstrated a speedup of nearly $5$ times for Fashion-MNIST, $10$ times for ImageNet, and $20$ times for the padded ImageNet dataset.
Figure 4: Speedup of pyECT over FastTopology across multiple cores of ECF computation on Fashion-MNIST, ImageNet, and a padded ImageNet
5 Discussion
In this work, we presented a vectorized framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension with no structural constraints. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of image datasets. The implementation is also incredibly concise as the core vectorized operations leverage PyTorch, a well-maintained and optimized library for tensor computation. The core WECT algorithm in the pyECT implementation is only about $25$ lines of code, which although we acknowledge it is not a perfect metric, it is a testament to the simplicity of the approach.
The pyECT package also contains a variety of tools for constructing and working with simplicial/cubical complexes. Our package includes fully vectorized methods for building weighted Freudenthal complexes and weighted cubical complexes from images. Additionally, pyECT has dedicated methods for the specific cases of computing ECFs of images and WECTs of weighted geometric simplicial complexes in $\mathbb{R}^{n}$. These two special cases are implemented as PyTorch neural network modules.
Another benefit of the approach presented in this work is that vectorized operations support automatic gradient computations. Potential future work includes exploring the use of this framework in topological machine learning applications and, in particular, incorporating topological transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
Acknowledgments
Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University. ChatGPT was used in a brain-storming and debugging capacity for this work.
Appendix A Examples of Tensor Operations
In this appendix, we provide a few examples to illustrate the tensor operations defined in Section 2.4. Consider the following tensors:
$T=\begin{bmatrix}1&2\\ 3&4\\ 5&6\end{bmatrix},\quad U=\begin{bmatrix}1\\ 2\end{bmatrix},\quad V=\begin{bmatrix}5\end{bmatrix}$
The CumulativeSum operation computes the cumulative sum of each row of the tensor. For the tensor $T$, we see
$\texttt{CumulativeSum}(T)=\begin{bmatrix}1&1+2\\ 3&3+4\\ 5&5+6\end{bmatrix}=\begin{bmatrix}1&3\\ 3&7\\ 5&11\end{bmatrix}$.
The ReduceMax operation computes the maximum value with respect to a specified axis. That is, for the tensor $T$ and axis $p=1$, we have
$\texttt{ReduceMax}(T,1)=\begin{bmatrix}\max(1,2)\\ \max(3,4)\\ \max(5,6)\end{bmatrix}=\begin{bmatrix}2\\ 4\\ 6\end{bmatrix}$.
Next, we consider the advanced indexing operation, using $U$ to index into $T$. Note that $T$ has shape $(3,2)$ and $U$ has shape $(2,1)$, so the resulting tensor has shape $(2,1,2)$. Then, for each $k\in[[2]]$ we have
$T[U]=\begin{bmatrix}T[U[0,0],k]\\ T[U[0,1],k]\end{bmatrix}=\begin{bmatrix}T[1,k]\\ T[2,k]\end{bmatrix}=\begin{bmatrix}\begin{bmatrix}3&4\end{bmatrix}\\ \begin{bmatrix}5&6\end{bmatrix}\end{bmatrix}$.
To demonstrate the ScatterAdd operation, we need to ensure dimensions align correctly, so consider the transpose $S=T^{T}$, which is
$S=\begin{bmatrix}1&3&5\\ 2&4&6\end{bmatrix}$,
Next, we compute $\texttt{ScatterAdd}(U,V)(S)$. To do so, we compute the difference tensor $D$, which is used to modify $S$ in place. The tensor $D$ has the same shape as $S$, which is $(2,3)$. Because $V$ has shape $(1)$, we note that $k=0$ is the only possible value of $k$ in the definition of $D$ given in Equation (4). As a result, we look for indices $i,j$ such that $U[i,0]=j$. Thus, we start with $D=\texttt{zeros}((2,3))$ and compute the values of $D$ by adding the following values:
  1. 1.
    For the first row ($i=0$), the index from $U$ is $U[0,0]=1$, and the value from $V$ is $5$. Thus, we add the value $5$ to the element at position $(i,U[i,k])=(0,U[0,0])=(0,1)$, so the first row of $D$ is $[0,5,0]$.
  2. 2.
    For the second row ($i=1$), the index from $U$ is $U[1,0]=2$, and the value from $V$ is $5$. Thus, we add the value $5$ to the element at position $(i,U[i,k])=(1,U[1,0])=(1,2)$, so the second row of $D$ is $[0,0,5]$.
Thus, the difference tensor is
$D=\begin{bmatrix}0&5&0\\ 0&0&5\end{bmatrix}$,
and we have
$\texttt{ScatterAdd}(U,V)(S)=S+D=\begin{bmatrix}1+0&3+5&5+0\\ 2+0&4+0&6+5\end{bmatrix}=\begin{bmatrix}1&8&5\\ 2&4&11\end{bmatrix}$.
Recall that the ScatterAdd operation modifies the input tensor in place, so the new value of $S$ is
$S=\begin{bmatrix}1&8&5\\ 2&4&11\end{bmatrix}$.