### v2

# Vectorized Computation of Euler Characteristic Functions and Transforms

Abstract
The weighted Euler characteristic transform (WECT) and Euler characteristic function (ECF) have proven to be useful tools in a variety of applications. However, current methods for computing these functions are either not optimized for speed or do not scale to higher-dimensional settings. In this work, we present a vectorized framework for computing such topological descriptors using tensor operations, which is highly optimized for GPU architectures and works in full generality across simplicial and cubical complexes of arbitrary dimension. Experimentally, the framework demonstrates significant speedups over existing methods when computing the WECT and ECF across a variety of two- and three-dimensional datasets. Computation of these transforms is implemented in a publicly available Python package called pyECT.
1 Introduction
The field of topological data analysis (TDA) utilizes invariants from topology to study the shape of data. Among the invariants employed are homology and the original topological invariant: the Euler characteristic [10]. In practice, a single invariant is not typically robust enough to capture the shape of data. For this reason, filtrations (i.e., nested sequences of shapes) are often constructed from data and topological invariants are then applied. This allows for significantly more information to be retained, leading to constructions such as persistence diagrams [9] and Euler characteristic functions [20].
Topological transforms [23] further extend the idea of applying topological invariants to filtrations by constructing a parameterized family of filtrations and then applying topological invariants. The starting point for topological transforms is typically a geometric simplicial or cubical complex $K$ in $\mathbb{R}^{n}$. For each direction vector $\mathbf{s}\in\mathbb{S}^{n-1}$, a directional filtration of $K$ is constructed and then a topological invariant is applied to that filtration. Applying the Euler characteristic (EC) to a single directional filtration results in the Euler characteristic function (ECF), while applying it to each directional filtration in $\mathbb{S}^{n-1}$ results in the Euler characteristic transform (ECT) [23, 15]. The ECT is a complete invariant: a geometric simplicial complex $K$ in $\mathbb{R}^{n}$ can be reconstructed solely from its ECT [6, 11]. In other words, no information is lost when passing from the simplicial complex $K$ to its ECT.
Useful variants of the ECF and ECT are the weighted Euler characteristic function (WECF) and the weighted Euler characteristic transform (WECT) for weighted complexes [12]. Much like the ECT, the WECT is a complete invariant for weighted geometric complexes [12]. Moreover, the WECT has been employed in a broad range of applications, including: in oncology for predicting survival rates of patients with glioblastoma multiforme brain tumors [5], in plant biology for classifying seed phenotypes [1], and in image classification tasks [3, 12]. Due to its popularity, numerous software packages already exist for computing the WECT and its variants, including those presented in [1, 5, 13, 14, 17, 18]. While many of these packages provide useful tools, they are either not optimized for speed or are limited in the types (or dimension) of complexes they can process.
In this work, we provide a fully vectorized framework for computing WECFs and WECTs. Our implementation applies in full generality to arbitrary geometric simplicial and cubical complexes in Euclidean spaces. The vectorized operations are highly optimized for fast computation on modern GPU hardware, providing speedups over existing implementations.
2 Background
In this section, we provide necessary foundations in computational topology and tensor computation to develop the main results. For simplicity, we restrict our attention to simplicial complexes in this paper; however, our results and code work for cubical complexes.
We use abstract simplicial complexes throughout this work. For a (abstract) simplicial complex $K$ and for any $i\in\{0,1,\ldots,\dim(K)\}$, we denote the set of $i$-simplices of $K$ by $K_{i}$ and treat $\sigma\in K$ as a set of vertices. Furthermore, we define $k=|K|$ and $k_{i}=|K_{i}|$.
2.1 Lower-Star Filtrations
Here, we provide the necessary background on filtrations.
Definition 1 (Filtrations and Filters).
Let $K$ be a simplicial complex and let $\operatorname{sub}(K)$ denote the set of subcomplexes of $K$. A filtration of $K$ is a function $F\colon\mathbb{R}\to\operatorname{sub}(K)$} such that for every $s\leq t\in\mathbb{R}$, we have a nesting of the respective complexes: $F(s)\subseteq F(t)$. [^1] This definition generalizes to any poset as the domain of $F$ rather than $\mathbb{R}$. However, the one-parameter setting is sufficient for our purposes.
A filter on $K$ is a function $f\colon K\to\mathbb{R}$ such that the function $F:\mathbb{R}\to\operatorname{sub}(K)$ defined by
$F(t):=\{\sigma\in K\mid f(\sigma)\leq t\}$ (1)
is a filtration.
Next, we introduce a class of filtrations that are the focus of this work.
Definition 2 (Lower-Star Filtrations).
Let $K$ be a simplicial complex with vertex set $K_{0}$. A vertex filter on $K$ is a function $f\colon K_{0}\to\mathbb{R}$}. For a vertex filter $f:K_{0}\to\mathbb{R}$, the lower-star filtration induced by $f$ is the function $F:\mathbb{R}\to\operatorname{sub}(K)$, defined as
$F(t):=\{\sigma\in K\mid\max_{v\in\sigma}f(v)\leq t\}.$ (2)
The main examples of lower-star filtrations used in this paper are directional filtrations. These are defined for any (abstract) simplicial complex $K$ with a function $\ell\colon K_{0}\to\mathbb{R}^{n}$. In practice, simplicial complexes are often naturally embedded in $\mathbb{R}^{n}$ and, in this setting, the function $\ell$ simply maps each vertex to its coordinates.
Definition 3 (Vertex Height Functions and Directional Filtrations).
Let $K$ be a simplicial complex with a map $\ell:K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The vertex height function in direction $\mathbf{s}$ is the map $h_{\mathbf{s}}:K_{0}\to\mathbb{R}$} defined as $h_{\mathbf{s}}(v):=\ell(v)\cdot\mathbf{s}.$ The directional filtration of $K$ in direction $\mathbf{s}$, denoted as $H_{\mathbf{s}}$, is the lower-star filtration of $K$ induced by $h_{\mathbf{s}}$.
2.2 The Weighted Euler Characteristic Transform
Here, we introduce the main focus of this paper: the WECT.
Definition 4 (Weighted Simplicial Complexes).
A weighted simplicial complex is a pair $(K,\omega)$, where $K$ is a simplicial complex and $\omega:K\to\mathbb{R}$ is a function (that we call a weight function).
Weighted simplicial complexes are often used to model various types of data. For example, triangular meshes are widely used to model three-dimensional surfaces, with weight functions capturing scalar fields on the surface. Another example comes from image data, where grayscale images are converted into weighted simplicial complexes via the Freudenthal triangulation [3, 12] with weights derived from the pixel intensities.
Originally introduced by Euler in 1758 for convex polyhedra [10], the Euler characteristic (EC) is a topological invariant that readily extends to simplicial complexes. We use an extension of the EC to weighted simplicial complexes: the weighted Euler characteristic.
Definition 5 (Weighted Euler Characteristic).
The weighted Euler characteristic (WEC) of a weighted simplicial complex $(K,\omega)$ is
$\chi(K,\omega):=\sum_{\sigma\in K}(-1)^{\dim{\sigma}}\omega(\sigma).$
Note that the Euler characteristic is a special case of the WEC, where the weight function is equal to the constant function $\mathbf{1}:K\to\mathbb{R}$. While the WEC is a remarkably useful invariant given its simplicity, there is limited information that a single number can capture about a simplicial complex. To incorporate more information, we apply the WEC to a filtration:
Definition 6 (Weighted Euler Characteristic Function).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex and let $F\colon\mathbb{R}\to\operatorname{sub}(K)$ be a filtration of $K$. The weighted Euler characteristic function [^2] Weighted Euler characteristic functions are also known as weighted Euler characteristic curves. Due to their general lack of continuity, we use the term function rather than curve. of $\mathbb{K}$ induced by $F$ is the function $\operatorname{WECF}_{F}:\mathbb{R}\to\mathbb{R}$ defined as
$\operatorname{WECF}_{F}(t):=\chi\big(F(t),\omega|_{F(t)}\big).$ (3)
When $F$ is induced from a (vertex) filter function $f$, we write $\operatorname{WECF}_{f}:=\operatorname{WECF}_{F}$.
Definition 7 (Directional WECF).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a function $\ell\colon K_{0}\to\mathbb{R}^{n}$ and let $\mathbf{s}\in\mathbb{S}^{n-1}$} be a direction vector. The directional WECF of $\mathbb{K}$ in direction $\mathbf{s}$ is the WECF of $\mathbb{K}$ induced by the height function $h_{\mathbf{s}}$.
Remark 8.
In image analysis tasks (and their three-dimensional analogs), there are two common choices of vertex filters considered. The first (implemented in [13, 24]) uses the pixel intensities as the vertex filter. It then computes the result by applying the (unweighted) Euler characteristic to the corresponding lower-star filtration. We refer to this simply as the ECF of an image. The second, which we refer to as the directional WECF of an image, uses the pixel intensities as weights and the height function $h_{\mathbf{s}}$ for a given direction vector $\mathbf{s}\in\mathbb{S}^{1}$ as the vertex filter. This method is implemented in [1, 22, 14].
Example 9.
An illustration of the directional ECF and directional WECF is presented in Figure 1. The full simplicial complex, including weighted vertices, is displayed in Figure 1(1(c)). The filtration is along direction $(1,0)$, and the corresponding ECF and WECF are included in Figure 1(1(d)) and Figure 1(1(e)), respectively. The ECF and WECF are the same until time 1 when two new vertices are introduced with weights of 0.5.
(a) Time=0
(b) Time=1
(c) Time=2
(d) ECF in direction (1,0)
(e) WECF in direction (1,0)
Figure 1: Filtration of a simplicial complex (1(a))–(1(c)) along direction (1,0), with vertex weights shown in green discs. The weight assigned to a simplex $\sigma$ is the maximum weight of its zero-simplex faces. The corresponding ECF (1(d)) and WECF (1(e)) are displayed in the; see Equation (3). At times 1 and 2, the ECF and WECF differ due to the WECF’s use of the weighted simplices.
Typically, there is no canonical direction vector to choose when computing directional WECFs. Hence, it is common to choose all directions resulting in the weighted Euler characteristic transform (WECT).
Definition 10 (Weighted Euler Characteristic Transform).
Let $\mathbb{K}=(K,\omega)$ be a weighted simplicial complex with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$. The weighted Euler characteristic transform of $\mathbb{K}$ is the function $\operatorname{WECT}_{\mathbb{K}}:\mathbb{S}^{n-1}\times\mathbb{R}\to\mathbb{R}$ defined as
$\operatorname{WECT}_{\mathbb{K}}(\mathbf{s},t):=\operatorname{WECF}_{h_{\mathbf{s}}}(t)=\chi\big(H_{\mathbf{s}}(t),\omega\big).$ (4)
2.3 Existing Frameworks for Computation of the ECF and (W)ECT
Due to its simplicity and ease of computation, the WECT and its variants are widely used in applied settings (resulting in many available software packages). Demeter [1] is a Python package that computes the directional ECF and ECT of axis-aligned cubical complexes. Note that it incorporates an optimization of bucket sorting the filtration values of each vertex. However, this package does not work for generic simplicial complexes, and only computes the special case of the unweighted ECT. Next, the SINATRA package [22] is an R package that computes the directional ECF and ECT through an algorithm similar to Demeter. However, unlike Demeter, SINATRA works for arbitrary, unweighted simplicial complexes. Note that as Demeter and SINATRA are not optimized for speed, they are not included in the experimental comparison. Third, EuCalc  [^3] https://github.com/HugoPasse/Eucalc [14] is a highly optimized package that uses a preprocessing step to compute critical points of the directional ECF. This allows for efficient computation of the exact directional ECF (as opposed to a vectorization of it). The package only accepts axis-aligned cubical complexes as input, but does generalize to weighted complexes. Finally, the DECT package [^4] https://github.com/aidos-lab/dect [17] is a Python package designed to compute the differentiable ECT (DECT) of simplicial complexes. While the DECT is a slightly different variant, the package also includes an implementation for computing the ECT of simplicial or cubical complexes. Although it does not allow for weighted complexes, this package is designed for GPU architectures making it an interesting point of comparison. Furthermore, [18] provides further optimizations for computing the DECT on CUDA GPU architectures, but is consequently a far less general framework.
The FastTopology framework [^5] https://github.com/zavalab/ML/tree/master/FastTopology [13] is a highly optimized implementation for computing the ECF of lower-star filtrations of two- and three-dimensional axis-aligned cubical complexes. While it is designed to be highly parallelizable on CPU architectures, it does not directly support vectorized computation (e.g., on a GPU) or the WECF. GPU ECC [24] is a GPU-based implementation of the ECF. Unlike the other existing packages, it is designed for GPU architectures. However, it is highly optimized for image data, so it only works for axis-aligned cubical complexes and does not generalize to weighted complexes. Additionally, it does not compute the directional ECF or ECT, and currently is only available as an implementation on Windows systems (which is unavailable in our computation environment).
The framework presented in this paper, pyECT, admits several advantages over these existing packages for computing the WECT and ECF. First, unlike all existing frameworks (except SINATRA), pyECT works in full generality across axis-aligned cubical complexes, arbitrary cubical complexes, and geometric simplicial complexes; see Table 1. Second, pyECT provides frameworks for computing both the WECFs of lower-star filtrations and WECTs of weighted geometric complexes. This generality is not available in any of these existing packages. Third, the vectorized operations used in pyECT are highly optimized for modern GPU hardware, which, in most cases, results in significant speedups over these existing packages. The GPU ECC package admits this advantage as well, but is far less general as it only accepts axis-aligned cubical complexes and does not compute the directional WECF or WECT. Finally, the simplicity of the vectorized approach makes it very advantageous for software implementation, especially in the context of well-maintained, abstracted tensor computation libraries like PyTorch [16, 2]. For example, the core WECT algorithm in the pyECT implementation is only about $25$ lines of code.
One minor disadvantage of pyECT is that it does not compute exact EC values at all heights of the directional ECF (which is available in EuCalc). While computing exact values retains more information and can be an advantage for reconstruction tasks, it leads to difficulties in vectorizing the (W)ECT. However, vectorizations of the WECT are in fact more desirable for many real-world applications, especially when used as input to distance-based methods or machine learning pipelines (such as those in [12, 23, 15, 1]). While it is possible to compute a vectorization from exact values using EuCalc, it requires an additional computation step that is linear in the size of the discretization.
Table 1: Comparison of acceptable input types of ECT packages.
Package Simplicial Complexes Cubical Complexes Axis-Aligned Cubical Complexes
pyECT yes yes yes
EuCalc no no yes
FastTopology no no yes
Demeter no no yes
SINATRA yes yes yes
2.4 Tensor Operations
Our algorithms almost exclusively use vectorized operations on tensors. These have two main advantages over non-vectorized approaches. First, the vectorized operations we use have highly optimized implementations in PyTorch [16, 2] that run in parallel on both CPUs and GPUs. Second, they support automatic gradient computations, facilitating the use of differentiable versions in our implementations in neural networks. Here, we review tensors and the vectorized operations used in our algorithm.
For any positive integer $m$, let $m]]:=\{0,1,\ldots,m-1\}$. For $\{m_{i}\}_{i=0}^{d-1}\subset\mathbb{Z}_{+}$, a $d$-dimensional tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ is a function
$T\colon[[m_{0}]]\times[[m_{1}]]\times\cdots\times[[m_{d-1}]]\to\mathbb{R}.$
We exclusively use one-dimensional tensors (i.e., vectors), two-dimensional tensors (i.e., matrices), and three-dimensional tensors. A specific dimension of a tensor is called an axis. We now introduce the tensor operations utilized in our algorithm (and refer the reader to Appendix A for examples).
  1. 1.
    Zeros: Given a shape $L$, $\texttt{zeros}(L)$ is the tensor with shape $L$, where every value is 0.
  2. 2.
    Element-wise operations: Any function $f:\mathbb{R}\to\mathbb{R}$ can be applied to tensors element-wise by post-composition. That is, if $T$ is a tensor, then $f(T)$ is the tensor with the same shape as $T$ and values
    $f(T)[i_{0},i_{1},\ldots,i_{d-1}]=f\big(T[i_{0},i_{1},\ldots,i_{d-1}]\big).$
  3. 3.
    Matrix multiplication: If $T$ and $S$ are tensors with shape $(m_{0},m_{1})$ and $(m_{1},m_{2})$ respectively, we use the notation TST*S to denote their matrix multiplication.
  4. 4.
    Cumulative sum: Let $T$ be a tensor of shape $(m_{0},m_{1})$. Then, $\texttt{CumulativeSum}(T)$ is the tensor with shape $(m_{0},m_{1})$ and values
    $\texttt{CumulativeSum}(T)[i,j]=\sum_{\begin{subarray}{c}k\in[[m_{1}]]\\ k\leq j\end{subarray}}T[i,k].$
    In this work, we use the cumulative sum on two-dimensional tensors over the second axis, $m_{1}]]$. However, the cumulative sum can be defined over any axis of a tensor.
  5. 5.
    Reduce maximum: Let $T$ be a tensor of shape $(m_{0},m_{1},\ldots,m_{d-1})$ and let $p\in[[d]]$. Then, $\texttt{ReduceMax}(T,p)$ is the tensor with shape $(m_{0},\ldots,m_{p-1},m_{p+1},\ldots,m_{d-1})$ and values
    $\texttt{ReduceMax}(T,p)[i_{0},\ldots,i_{p-1},i_{p+1},\ldots i_{d-1}]=\max_{j\in[[m_{p}]]}T[i_{0},\ldots,i_{p-1},j,i_{p+1},\ldots i_{d-1}].$
  6. 6.
    Advanced indexing: Let $T$ be a tensor with shape $(m_{2},m_{3})$, and let $I$ be a tensor with shape $(m_{0},m_{1})$ and values in $m_{2}]]$. The tensor $T[I]$ has shape $(m_{0},m_{1},m_{3})$ and values
    $T[I][i,j,k]=T\big[I[i,j],k\big].$
  7. 7.
    Scatter add: Let $S$ be a tensor with shape $(m_{0},m_{1})$, let $I$ be a tensor with shape $(m_{0},m_{2})$ and values in $m_{1}]]$, and let $T$ be a tensor with shape $(m_{2})$. The function $\texttt{ScatterAdd}(I,T)$ updates $S$ in-place by adding to it, the difference tensor $D$ with shape $(m_{0},m_{1})$ and values
    $D[i,j]=\sum_{\begin{subarray}{c}k\in[[m_{2}]]\\ I[i,k]=j\end{subarray}}T[k].$ (5)
    That is, for a tensor $S$, we have $\texttt{ScatterAdd}(I,T)(S)=S+D,$ where ++ is taken element-wise and $S$ is updated in place. In practice, the tensor $D$ is not explicitly built.
3 Computation of WECFs
In this section, we present a new algorithm for computing the WECFs of a weighted simplicial complex with respect to a finite set of vertex filters. Computation of the WECT of a weighted geometric simplicial complex is a special case of Algorithm 1, and corresponds to the case where the vertex filters are height functions; see Example 11.
We fix a weighted simplicial complex $\mathbb{K}=(K,\omega)$ and a set of vertex filters $\{f_{0},f_{1},\ldots,f_{m-1}\}$ for the entirety of this section. To simplify notation, we extend each $f_{p}:K_{0}\to\mathbb{R}$ to $K$ by defining $f_{p}^{\max}:K\to\mathbb{R}$ as
$f_{p}^{\max}(\sigma)=\max_{v\in\sigma}f_{p}(v).$
We fix enumerations $K_{0}=\{v_{0},v_{1},\ldots,v_{k_{0}}\}$ and $K_{i}=\{\sigma_{0}^{i},\sigma_{1}^{i},\ldots,\sigma_{k_{i}}^{i}\}$ for each $i\geq 1$.
Our algorithm uses the following inputs
  • Complex: A list of tuples:
    • $\texttt{Complex}[0]=(\texttt{Filters},\texttt{VertexWeights})$
        Filters is a tensor with shape $(k_{0},m)$ and values $\texttt{Filters}[a,p]=f_{p}(v_{a}).$
        VertexWeights is a tensor with shape $(k_{0})$ and values $\texttt{VertexWeights}[a]=\omega(v_{a}).$
    • for $i\geq 1$: $\texttt{Complex}[i]=(i\texttt{-SimplexVertices},i\texttt{-SimplexWeights})$
        $i\texttt{-SimplexVertices}$ is a tensor with shape $(k_{i},i+1)$ and values in $k_{0}]]$. The rows of this tensor correspond to the vertex sets of each $i$-simplex.
        $i\texttt{-SimplexWeights}$ is a tensor with shape $(k_{i})$ and values $i\texttt{-SimplexWeights}[b]=\omega(\sigma_{b}^{i}).$
  • NumVals: The number of values to sample the filter functions over.
In order to discretize the filtration height parameter, define MaxVal as
$\texttt{MaxVal}=\max_{\begin{subarray}{c}p\in[[m]]\\ v\in K_{0}\end{subarray}}|f_{p}(v)|.$
We discretize $\texttt{MaxVal},\texttt{MaxVal}]$ by evenly sampling NumVals points. That is, we define the map
$\beta\colon[[\texttt{NumVals}]]\to[-\texttt{MaxVal},\texttt{MaxVal}]$
by
$\beta(q)=q\frac{2*\texttt{MaxVal}}{\texttt{NumVals}-1}-\texttt{MaxVal}.$
The function $\beta$ has a left inverse $\alpha\colon[-\texttt{MaxVal},\texttt{MaxVal}]\to[[\texttt{NumVals}]]$ defined as
$\alpha(t)=\bigg\lceil\frac{(\texttt{NumVals}-1)(\texttt{MaxVal}+t)}{2*\texttt{MaxVal}}\bigg\rceil.$ (6)
The maps $\alpha$ and $\beta$ are order-preserving and satisfy
$\alpha(t)\leq q\iff t\leq\beta(q)$ (7)
for all $t\in[-\texttt{MaxVal},\texttt{MaxVal}]$ and $q\in[[\texttt{NumVals}]]$. [^6] In other words, $\alpha$ and $\beta$ form a Galois connection.
[…]
0:Complex, NumVals
0:WECFs: a tensor of shape $(m,\texttt{NumVals})$ and values $\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}(\beta(q))$.
[…]
5:for $i=1$ to $\dim(K)$ do
6:  SimplexVertices, SimplexWeights $\leftarrow\texttt{Complex}[i]$
7:  SimpIndices $\leftarrow\texttt{VIndices}[\texttt{SimplexVertices}]$
8:  $\texttt{MaxSimpIndices}\leftarrow\texttt{ReduceMax}(\texttt{SimpIndices},1)$
9:  $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})(\texttt{diffWECFs})$
10:end for
[…]
12:return WECFs
Algorithm 1 ultimately produces a tensor WECFs with shape $(m,\texttt{NumVals})$ and values
$\texttt{WECFs}[p,q]=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$
The algorithm begins by initializing a zero-tensor diffWECFs with shape $(m,\texttt{NumVals})$. Next, the vertex filter values are mapped to indices in $\texttt{NumVals}]]$ by applying the map $\alpha$ elementwise to the tensor Filters. This gives the tensor VIndices with values
$\texttt{VIndices}[a,p]=\alpha\big(f_{p}(v_{a})\big).$ (8)
The operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is then applied to diffWECFs, resulting in adding to diffWECFs, the tensor $D_{0}$ with shape $(m,\texttt{NumVals})$ and values
$D_{0}[p,q]=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]=q\end{subarray}}\texttt{VertexWeights}[a].$
Next, inside the for-loop, fix an $i\in\{1,2,\ldots\dim(K)\}$. The tensor SimplexVertices has rows indexed by $K_{i}$ with the entries in each row corresponding to the vertex sets of simplices in $K_{i}$. Therefore, the tensor MaxSimpIndices has shape $(k_{i},m)$ and values
$\texttt{MaxSimpIndices}[b,p]$ $=\max_{j\in[[i+1]]}\texttt{VIndices}\big[\texttt{SimplexVertices}[b,j],p\big]$
$=\max_{v\in\sigma_{b}^{i}}\alpha\big(f_{p}(v)\big)$
$=\alpha\big(\max_{v\in\sigma_{b}^{i}}f_{p}(v)\big)\hskip 42.67912pt\text{as }\alpha\text{ is order-preserving}$
$=\alpha\big(f_{p}^{\max}(\sigma_{b}^{i})\big).$ (9)
The operation $\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$ updates diffWECFs by adding to it, the tensor $D_{i}$ with shape $(m,\texttt{NumVals})$ and values
$D_{i}[p,q]=\sum_{\begin{subarray}{c}b\in[[k_{i}]]\\ \texttt{MaxSimpIndices}[b,p]=q\end{subarray}}(-1)^{i}\texttt{SimplexWeights}[b].$
After the conclusion of the for loop, we have $\texttt{diffWECFs}=\sum_{i=0}^{\dim(K)}D_{i}.$ Finally, a cumulative sum is applied to diffWECFs to obtain the tensor WECFs. From the fact that the cumulative sum is a linear operator, we have
$\texttt{WECFs}=\sum_{i=0}^{\dim(K)}\texttt{CumulativeSum}(D_{i}).$
The cumulative sum of $D_{0}$ has values
$\texttt{CumulativeSum}(D_{0})[p,q]$ $=\sum_{q^{\prime}\leq q}\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]=q^{\prime}\end{subarray}}\texttt{VertexWeights}[a]$
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \texttt{VIndices}[a,p]\leq q\end{subarray}}\texttt{VertexWeights}[a]$
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ \alpha(f_{p}(v_{a}))\leq q\end{subarray}}\texttt{VertexWeights}[a]$ by Equation (8)
$=\sum_{\begin{subarray}{c}a\in[[k_{0}]]\\ f_{p}(v_{a})\leq\beta(q)\end{subarray}}\texttt{VertexWeights}[a]$ by Equation (7)
$=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v).$ (10)
Likewise, for any $i\in\{1,\ldots,\dim(K)\}$, Equation (3) together with the same argument used to show Equation (3) gives
$\texttt{CumulativeSum}(D_{i})[p,q]=\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma).$
Therefore, the tensor WECFs has values
$\texttt{WECFs}[p,q]$ $=\sum_{\begin{subarray}{c}v\in K_{0}\\ f_{p}(v)\leq\beta(q)\end{subarray}}\omega(v)+\sum_{i=1}^{\dim(K)}\sum_{\begin{subarray}{c}\sigma\in K_{i}\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
$=\sum_{\begin{subarray}{c}\sigma\in K\\ f_{p}^{\max}(\sigma)\leq\beta(q)\end{subarray}}\omega(\sigma)$
$=\operatorname{WECF}_{f_{p}}\big(\beta(q)\big).$
Example 11.
Computing a discretization of the WECT of a weighted simplicial complex $\mathbb{K}=(K,\omega)$} with a map $\ell\colon K_{0}\to\mathbb{R}^{n}$ is a special case of Algorithm 1 as follows. Consider a tensor $D$, with shape $(d,n)$, of direction vectors over which to sample the WECT. For any $p\in[[d]]$, let $\mathbf{s}_{p}\in\mathbb{R}^{n}$ be the $p$-th row of $D$. Let $V$ be the tensor, with shape $(k_{0},n)$, of the vertex coordinates of $K$. Then, the tensor $V*D^{T}$ has values
$(V*D^{T})[a,p]=h_{\mathbf{s}_{p}}(v_{a}).$
Therefore, using the tensor $V*D^{T}$ as the Filters tensor in Algorithm 1 produces a discretization of the WECT over the finite sample of directions in $D$.
The algorithm achieves a theoretically optimal final time and memory complexity of $\Theta(k)$, where $k$ is the number of simplices in $K$. We refer the reader to Appendix B for the full complexity analysis of Algorithm 1.
4 Experiments
In order to demonstrate the efficacy of our implementation, we evaluate and compare the runtime of Algorithm 1 with existing methods; see Section 2.3 on a variety of datasets. [^7] The implementation of pyECT can be accessed here: https://github.com/compTAG/pyECT. The source code for reproducing experiments (with instructions) is available here: https://github.com/compTAG/pyECT-experiments.
4.1 Experimental Design
Experiments were performed on a single node with a single CPU or a single GPU (unless otherwise specified). EPYC Rome CPUs and NVIDIA A100 GPUs were used. Runtimes for CPU experiments were recorded using the time Python library using the performance counter for benchmarking. For the GPU experiments, runtimes were recorded using PyTorch’s torch.cuda.Event API to prevent inaccurate timing from asynchronous GPU operations. The methods were evaluated on the following two- and three-dimensional datasets:
  • Fashion-MNIST [25]. Fashion-MNIST is a dataset consisting of $28\times 28$ grayscale images of clothing items. Its training partition consists of $60000$ images with equal representation across ten classes ($6000$ images per class). In our experiments, we use a random subset of $1000$ images sampled uniformly without replacement from the training partition, which are then converted to cubical complexes.
  • ImageNet [7, 19]. The ImageNet Sample Image dataset [19], which consists of $1000$ samples from the Large Scale Visual Recognition Challenge (ILSVRC) 2012-2017 subset of the ImageNet dataset [7] was used. All images in the sample were converted to grayscale using the ITU-R 601-2 luma transform within the Pillow Python library [4]. Then, the images were padded to the maximum dimensions of all images in the dataset, resulting in each image having uniform size of $1632\times 2400$ pixels. The padded pixels were assigned a pixel intensity value of $1.0$ (on a zero to one scale), and the images were converted to cubical complexes.
  • 3DCCs. A synthetic dataset of $1000$ three-dimensional cubical complexes was generated. Each complex has dimension $100\times 100\times 100$, with pixel intensities drawn i.i.d. from the uniform distribution on $[0,1]$.
  • Stanford Bunny and Armadillo [21]. The Stanford 3D Scanning Repository provides triangulated mesh models of three-dimensional objects. We use the Bunny and Armadillo models. The meshes were converted to (unweighted) simplicial complexes after preprocessing to de-duplicate vertices, edges, and faces. In the resulting simplicial complexes, the Bunny dataset consists of $2503$ vertices, $7473$ edges, and $4968$ faces. The Armadillo is significantly larger, consisting of $172974$ vertices, $518916$ edges, and $345944$ faces.
4.2 Results
We now present the experimental results, measured by the average speedup factor of our method over the existing methods. The average speedup factor is defined as the runtime of the baseline method (e.g., EuCalc, DECT, or FastTopology) divided by the runtime of the proposed method, pyECT, averaged across all images in the dataset.
WECT Results
In this section, we present results on the average speedup of pyECT over EuCalc and DECT when computing the WECT of the Fashion-MNIST and ImageNet datasets. As EuCalc does not provide a GPU implementation, all results for EuCalc are measured on the CPU. Also, as EuCalc does not natively provide a vectorized result, we provide results both including and excluding the vectorization time.
The results for Fashion-MNIST are shown in Figure 2. Across all settings, pyECT demonstrates speedups over both EuCalc and DECT. When comparing against EuCalc (Figure 2(2(a))–(2(d))), pyECT demonstrates a speedup in both CPU and GPU implementations—even without the vectorization step. However, the GPU implementation of pyECT significantly outperforms the CPU implementation as the number of directions increases, which is especially pronounced when including the vectorization time for EuCalc. As expected, when increasing the number of values used in the filtration, the speedup factor of pyECT over EuCalc increases drastically. When comparing against DECT (subfigures (2(e)) and (2(f))), pyECT also demonstrates significant speedups. While the computation speed is comparable for low numbers of directions and filtration heights, pyECT significantly outperforms DECT for larger choices of parameters. Also, as DECT is optimized for GPU architectures, the average speedup on GPU architectures is less pronounced than that of other comparisons. However, when constrained to CPU architectures, pyECT is much faster, indicating another advantage of pyECT’s general applicability for all scenarios.
(a) pyECT (CPU) vs. EuCalc
(b) pyECT (GPU) vs. EuCalc
(c) pyECT (CPU) vs. EuCalc (vectorized)
(d) pyECT (GPU) vs. EuCalc (vectorized)
(e) pyECT (CPU) vs. DECT (CPU)
(f) pyECT (GPU) vs. DECT (GPU)
Figure 2: A comparison for computing on the Fashion-MNIST dataset for WECT computation. Figures (2(a)) and (2(b)) compare pyECT with EuCalc excluding the vectorization time, (2(c)) and (2(d)) compare pyECT with EuCalc including the vectorization time, and (2(e)) and (2(f)) compare pyECT with DECT. CPU implementations of pyECT are shown on the left and GPU implementations are shown on the right. Observe that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights.
The average speedup results on the ImageNet are shown in Figure 3. When comparing against EuCalc (both with and without the vectorization step, in Figures 3(a)3(d)), pyECT demonstrates a significant speedup in both CPU and GPU implementations. This speedup is especially pronounced for the GPU implementation of pyECT, as it is much more stable to large numbers of directions and filtration height values. Due to the large size of the complexes in this dataset, the CPU speedup of pyECT is less pronounced than on the Fashion-MNIST dataset (especially for large numbers of directions and heights). Furthermore, the vectorization time of EuCalc is less impactful on this dataset, as the vectorization step is less significant compared to the overall runtime for the large complexes. In the GPU implementation of pyECT, we observe a trend where the average speedup increases as the number of height filtration values increases. While this behavior appears counterintuitive, in reality the vectorized code encounters less efficient memory access patterns at lower numbers of heights, which leads to a smaller advantage over EuCalc with the smaller tensor values. For a technical description of memory access patterns specific to the GPU hardware used in these experiments, we refer the reader to [8]. In the comparison against DECT (Figures 3(e) and 3(f)), pyECT again demonstrates significant speedups. In fact, due to the large size of the complexes in this dataset, DECT was unable to compute the WECT for large numbers of directions and height values (due to memory constraints on the GPU). This results in the appearance of missing values in Figures 3(e) and 3(f). Note, however, that pyECT is still able to compute the WECT in these scenarios, demonstrating its memory efficiency. Furthermore, in the cases where the WECT could be computed by DECT, the pyECT implementation outperforms DECT for larger numbers of directions and height filtration values.
(a) pyECT (CPU) vs. EuCalc
(b) pyECT (GPU) vs. EuCalc
(c) pyECT (CPU) vs. EuCalc (vectorized)
(d) pyECT (GPU) vs. EuCalc (vectorized)
(e) pyECT (CPU) vs. DECT (CPU)
(f) pyECT (GPU) vs. DECT (GPU)
Figure 3: A comparison for computing on the ImageNet dataset for WECT computation. Figures (3(a)) and (3(b)) compare pyECT with EuCalc excluding the vectorization time, (3(c)) and (3(d)) compare pyECT with EuCalc including the vectorization time, and (3(e)) and (3(f)) compare pyECT with DECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms the baseline methods. This behavior is especially pronounced when using the GPU architecture with large numbers of directions and filtration heights. Note that DECT was unable to compute the WECT for large numbers of directions and heights due to memory constraints, resulting in missing values in (3(e)) and (3(f)). However, pyECT was able to compute the WECT in these scenarios.
Speedup results for the comparison of pyECT against DECT on the three-dimensional mesh datasets are presented in Figure 4. Note that experiments were run 10 times on each mesh dataset to generate the average and standard deviation values shown in the plots. For large choices of directions and filtration heights in both datasets, DECT was unable to compute the ECT due to memory constraints. While this results in missing values in all subplots, from the cases where DECT was able to compute the ECT, pyECT outperforms DECT (especially for larger choices of parameters). The memory issues are more pronounced in the Armadillo dataset (due to its larger size). Note that while pyECT demonstrates significant speedups in the GPU implementations, its advantage is even more pronounced in the CPU implementations, highlighting the versatility of pyECT across different hardware architectures.
(a) Bunny: pyECT (CPU) vs. DECT (CPU)
(b) Bunny: pyECT (GPU) vs. DECT (GPU)
(c) Armadillo: pyECT (CPU) vs. DECT (CPU)
(d) Armadillo: pyECT (GPU) vs. DECT (GPU)
Figure 4: A comparison of pyECT against DECT for computing the ECT on the Stanford Bunny and Armadillo meshes. Note that vertical axes are not consistent across subfigures. In many cases, DECT was unable to compute the ECT due to memory constraints (leading to missing values in the plots). However, pyECT was able to compute the ECT in all scenarios, and especially outperforms DECT for large numbers of directions and filtration heights.
ECF Results
In this section, we present the speedup of pyECT over FastTopology when computing the ECF of the Fashion-MNIST, ImageNet, and 3DCCs datasets. Due to the consistent computation graph structure of the ECF algorithm, pyECT is able leverage the PyTorch TorchDynamo compilation framework. These compiled runtimes are also measured and compared with FastTopology. As FastTopology is parallelizable across multiple CPU cores, we include results for FastTopology using $1$, 22, 44, 88, $12$, and $16$ CPU cores. Note that FastTopology does not provide a GPU implementation. The results are presented in Figure 5. We observe that pyECT provides a significant speedup over FastTopology in nearly all cases. When constrained to CPU architectures, pyECT has a less pronounced speedup, but is still comparable in speed on a single CPU core when compared with the fully parallelized FastTopology implementation. When utilizing GPU architectures, pyECT demonstrates a significant speedup over the fully parallelized FastTopology, especially on the three-dimensional 3DCCs dataset. The compilation strategy further increases the speed of pyECT on GPU architectures, but adds unrewarded overhead on CPU architectures.
(a) pyECT (CPU) vs. FastTopology
(b) pyECT (GPU) vs. FastTopology
(c) pyECT (compiled, CPU) vs. FastTopology
(d) pyECT (compiled, GPU) vs. FastTopology
Figure 5: A comparison of average speedup of pyECT over FastTopology when computing the ECF of multiple datasets. The comparison evaluates both uncompiled and compiled versions of pyECT. Note that vertical axes are not consistent across subfigures. In nearly all cases, pyECT outperforms FastTopology. This behavior is especially pronounced when using the GPU architecture on the three-dimensional 3DCCs dataset.
5 Discussion
In this work, we present a vectorized framework for computing WECFs (including ECFs and the WECT). This framework allows for efficient computation of these functions using tensor operations, which are well-suited for modern GPU hardware. Furthermore, our method works in full generality for weighted simplicial complexes (or cubical complexes) of arbitrary dimension with no structural constraints. Our framework is implemented in a publicly available Python package called pyECT. Experimentally, pyECT demonstrates a significant speedup over existing methods across a variety of image datasets. The implementation is also incredibly concise as the core vectorized operations leverage PyTorch, a well-maintained and optimized library for tensor computation. The core WECT algorithm in the pyECT implementation is only about $25$ lines of code, which we acknowledge is not a perfect metric, but is a testament to the simplicity of the approach.
The pyECT package also contains a variety of tools for constructing and working with simplicial/cubical complexes. Our package includes fully vectorized methods for building weighted Freudenthal complexes and weighted cubical complexes from images. Additionally, pyECT has dedicated methods for the specific cases of computing ECFs of images and WECTs of weighted geometric simplicial complexes in $\mathbb{R}^{n}$. These two special cases are implemented as PyTorch neural network modules.
Another benefit of the approach presented in this work is that vectorized operations support automatic gradient computations. Potential future work includes exploring the use of this framework in differentiable forms of topological transforms (as in [17]) and, in particular, incorporating these transforms into neural network architectures. Additionally, we believe that this work will encourage further investigation into vectorized methods for computing other topological functions and transforms.
Acknowledgements
We thank Jack Ruder and Jacob Sriraman for their valuable assistance in developing and refining portions of the methodology and code used in this work. Computational efforts were performed on the Tempest High Performance Computing System, operated and supported by University Information Technology Research Cyberinfrastructure (RRID:SCR_026229) at Montana State University.
Appendix A Examples of Tensor Operations
In this appendix, we provide a few examples to illustrate the tensor operations defined in Section 2.4. Consider the following tensors:
$T=\begin{bmatrix}1&2\\ 3&4\\ 5&6\end{bmatrix},\quad U=\begin{bmatrix}1\\ 2\end{bmatrix},\quad V=\begin{bmatrix}5\end{bmatrix}$
The CumulativeSum operation computes the cumulative sum of each row of the tensor. For the tensor $T$, we see
$\texttt{CumulativeSum}(T)=\begin{bmatrix}1&1+2\\ 3&3+4\\ 5&5+6\end{bmatrix}=\begin{bmatrix}1&3\\ 3&7\\ 5&11\end{bmatrix}$.
The ReduceMax operation computes the maximum value with respect to a specified axis. That is, for the tensor $T$ and axis $p=1$, we have
$\texttt{ReduceMax}(T,1)=\begin{bmatrix}\max(1,2)\\ \max(3,4)\\ \max(5,6)\end{bmatrix}=\begin{bmatrix}2\\ 4\\ 6\end{bmatrix}$.
Next, we consider the advanced indexing operation, using $U$ to index into $T$. Note that $T$ has shape $(3,2)$ and $U$ has shape $(2,1)$, so the resulting tensor has shape $(2,1,2)$. Then, for each $k\in[[2]]$ we have
$T[U]=\begin{bmatrix}T[U[0,0],k]\\ T[U[0,1],k]\end{bmatrix}=\begin{bmatrix}T[1,k]\\ T[2,k]\end{bmatrix}=\begin{bmatrix}\begin{bmatrix}3&4\end{bmatrix}\\ \begin{bmatrix}5&6\end{bmatrix}\end{bmatrix}$.
To demonstrate the ScatterAdd operation, we need to ensure dimensions align correctly, so consider the transpose $S=T^{T}$, which is
$S=\begin{bmatrix}1&3&5\\ 2&4&6\end{bmatrix}$,
Next, we compute $\texttt{ScatterAdd}(U,V)(S)$. To do so, we compute the difference tensor $D$, which is used to modify $S$ in place. The tensor $D$ has the same shape as $S$, which is $(2,3)$. Because $V$ has shape $(1)$, we note that $k=0$ is the only possible value of $k$ in the definition of $D$ given in Equation (5). As a result, we look for indices $i,j$ such that $U[i,0]=j$. Thus, we start with $D=\texttt{zeros}((2,3))$ and compute the values of $D$ by adding the following values:
  1. 1.
    For the first row ($i=0$), the index from $U$ is $U[0,0]=1$, and the value from $V$ is $5$. Thus, we add the value $5$ to the element at position $(i,U[i,k])=(0,U[0,0])=(0,1)$, so the first row of $D$ is $[0,5,0]$.
  2. 2.
    For the second row ($i=1$), the index from $U$ is $U[1,0]=2$, and the value from $V$ is $5$. Thus, we add the value $5$ to the element at position $(i,U[i,k])=(1,U[1,0])=(1,2)$, so the second row of $D$ is $[0,0,5]$.
Thus, the difference tensor is
$D=\begin{bmatrix}0&5&0\\ 0&0&5\end{bmatrix}$,
and we have
$\texttt{ScatterAdd}(U,V)(S)=S+D=\begin{bmatrix}1+0&3+5&5+0\\ 2+0&4+0&6+5\end{bmatrix}=\begin{bmatrix}1&8&5\\ 2&4&11\end{bmatrix}$.
Recall that the ScatterAdd operation modifies the input tensor in place, so the new value of $S$ is
$S=\begin{bmatrix}1&8&5\\ 2&4&11\end{bmatrix}$.
Appendix B Complexity Analysis
Here, we analyze the asymptotic complexity of Algorithm 1. Assume that the ambient dimension $n$ is fixed. Let $k$ be the total number of simplices of $K$, let $m$ be the number of vertex filters, and let NumVals be the number of values to sample the vertex filters over.
The function ComputeWECFs has inputs Complex and NumVals requiring $\Theta(km)$ memory usage. The algorithm first initializes a zero tensor diffWECFs of shape $(m,\texttt{NumVals})$, requiring $\Theta(m\texttt{NumVals})$ memory. The function $\alpha$ (see Equation (6)) is then applied elementwise to Filters, taking $\Theta(km)$ time.
Next, the operation $\texttt{ScatterAdd}(\texttt{VIndices}^{T},\texttt{VertexWeights})$ is applied to diffWECFs. ScatterAdd adds a value from VertexWeights to diffWECFs for each entry of $\texttt{VIndices}^{T}$, resulting in a time complexity of $\Theta(km)$ and no additional memory usage as the operation is performed in place.
Now, for a fixed dimension $i\in\{1,2,\ldots n\}$, the tensor
$\texttt{SimpIndices}=\texttt{VIndices}[\texttt{SimplexVertices}]$
with shape $(k_{i},i+1,m)$ is constructed, taking $\Theta(km)$ time and memory. This tensor is then reduced along the second axis by taking the maximum values, creating the tensor MaxSimpIndices with shape $(k_{i},m)$ in $\Theta(km)$ time and memory. The tensor diffWECFs is then updated by applying the map
$\texttt{ScatterAdd}(\texttt{MaxSimpIndices}^{T},(-1)^{i}*\texttt{SimplexWeights})$
which takes $\Theta(km)$ time and memory.
Thus, we have a total time and memory complexity of $\Theta(nkm)=\Theta(km)$ after conclusion of the for-loop. Finally, a cumulative sum is applied to diffWECFs, taking $\Theta(m\texttt{NumVals})$ time and using no additional memory. This results in $\Theta(m(k+\texttt{NumVals}))$ for Algorithm 1 as the final time and memory complexity. In practice, $m$ and NumVals are typically held constant. Indeed, this is necessary to achieve consistent vectorizations of data. Holding $m$ and NumVals constant results in a theoretically optimal final time and memory complexity of $\Theta(k)$ for Algorithm 1.